### v2

@@ Abstract
For $p\in[1,\infty]$, we define a smooth manifold structure on the set $AC_{L^{p}}([a,b],N)$ of absolutely continuous functions $\gamma\colon[a,b]\to N$ with $L^{p}$-derivatives for all real numbers $a<b$ and each smooth manifold $N$ modeled on a sequentially complete locally convex topological vector space, such that $N$ admits a local addition. Smoothness of natural mappings between spaces of absolutely continuous functions is discussed, like superposition operators $AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2})$, $\eta\mapsto f\circ\eta$, for a smooth map $f\colon N_{1}\to N_{2}$. For $1\leq p<\infty$ and $r\in\mathbb{N}$ we show that the right half-Lie groups $\text{Diff}_{K}^{\text{ }r}(\mathbb{R}^{n})$ and $\text{Diff}^{\text{ }r}(M)$ are $L^{p}$-semiregular. Here $K$ is a compact subset of $\mathbb{R}^{n}$ and $M$ is a compact smooth manifold. An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\operatorname{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$. For the preceding examples, the evolution map $\operatorname{Evol}$ is continuous.
@@ 1 Introduction
Manifolds of mappings play an important role in various branches of global analysis, infinite-dimensional geometry and Lie theory. For a compact smooth manifold $M$, a finite-dimensional manifold $N$ and a non-negative integer $r$, a smooth manifold structure on the set $C^{r}(M,N)$ of $C^{r}$-functions $\phi\colon M\to N$ was first obtained by Eells (see [6] and the references therein). Later, manifold structures were also obtained on $C^{\infty}(M,N)$, and for infinite-dimensional manifolds $N$ as long as they admit a local addition, a concept recalled in Definition 1.1 (see [19, 15] and the references therein, also [2]). For a Hilbert manifold $M$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [7] and [14], cf. also [27]). For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [16] which is similar to Palais’ use of Banach section functors [25]. A Lie group structure (and hence a smooth manifold structure) on $AC_{L^{p}}([0,1],G)$ was obtained in [23] for each Lie group $G$ modeled on a sequentially complete locally convex space (as in [21]), generalizing the case of Fréchet–Lie groups treated in [10].
@@
defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $x\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
@@
has open image and is a $C^{\infty}$-diffeomorphism onto its image $\Omega^{\prime}$.
Throughout the following, $a<b$ are real numbers and $p\in[1,\infty]$. For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.11 and 2.21, respectively. For $\eta\in AC_{L^{p}}([a,b],N)$, the pointwise operations make
@@
${\mathcal{V}}_{\eta}\colon:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
@@
${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(t,\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
@@
If, in the situation of Definition 1.1, $N$ is $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition. In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{\eta}$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 4.6). If $N_{j}$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex topological $\mathbb{K}$-vector space such that $N_{j}$ admits a $\mathbb{K}$-analytic local addition for $j\in\{1,2\}$, then the map $AC_{L^{p}}([a,b],f)$ described in Theorem 1.3 is $\mathbb{K}$-analytic for all $p\in[1,\infty]$ (see Corollary 4.9).
@@
is a smooth map and a right action of $G$ in $TG$. The following concept strengthens “regularity” in the sense of Milnor [21].
@@
If $G$ is modeled on a Fréchet space, (1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure. In the general case, $\eta$ is required to be a Carathéodory solution to (1), i.e., it solves the corresponding integral equation piecewise in local charts. We mention that a Lie group $G$ is $L^{p}$-regular if and only if $G$ is $L^{p}$-semiregular and $\operatorname{Evol}$ is smooth as map
@@
(see [23]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [11, Theorem E]).
Now considetr a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$. Thus $G$ is a group, endowed with a smooth manifold structure modeled on $E$ which makes $G$ a topological group and turns the right translation $\rho_{g}\colon G\to G$ into a smooth mapping for each $g\in G$ (cf. [5, 18]). Let us use notation as in the case of Lie groups.
@@
$\dot{y}(t)=\gamma(t).y(t)$
satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [13]) and the initial value problem (1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
The Lie group $\text{Diff}(M)$ of $C^{\infty}$-diffeomorphisms of a compact smooth manifold $M$ without boundary is known to be $L^{1}$-regular, and also the Lie group $\text{Diff}_{K}(\mathbb{R}^{n})$ of all $C^{\infty}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\phi(x)=x$ for all $x\in\mathbb{R}^{n}\setminus K$, for each compact subset $K\subseteq\mathbb{R}^{n}$ (see [10]). For each positive integer $r$, the following analogues are obtained (see Theorems  7.5 and 6.3):
@@
If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the result remains valid.
@@
Let $1\leq p<\infty$. Let $G$ be a right half-Lie group modeled in a Fréchet space $E$ which admits local addition. Let $G$ be $L^{p}$-semiregular with continuous evolution map
@@
is continuous.
If we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
@@ 2 Preliminaries
exists for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$. Moreover, we assume that the mappings
@@
are continuous. Then the map $f$ is called a $C^{r,s}$-map.
@@
$\mu:\mathcal{B}(X)\to[0,\infty]$ on its $\sigma$-algebra of Borel sets and let $Y$ be a topological space. A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$ there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
@@
Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue measure on $\mathbb{R}$. We define the set $\mathcal{L}^{p}([a,b],E)$ as the set of all Lusin $\lambda$-measurable functions $\gamma:[a,b]\to E$ such that for each continuous seminorm $q$ on $E$ we have
@@
$\lVert\gamma\lVert_{\mathcal{L}^{q}}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
@@
We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ with compact and metrizable image, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$. Thus
@@
Let $E$ be a Frechet space and $p\in[0,1]$. If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [10]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([24, Proposition 2.10]).
@@
Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be a Lusin $\lambda$-measurable function such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$. An element $z\in E$ is called the weak integral of $\gamma$ if
$\forall\alpha\in E^{\prime}:\quad\alpha(z)=\int_{a}^{b}(\alpha\circ\gamma)(s)ds.$
@@
The following lemma [24, Proposition 2.26] allow us to define absolute continuous functions with vector values.
@@
We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$. Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique, the map
$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$
is an isomorphism of vector spaces. We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space, which is independent of the choice of $t_{0}$.
The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$ [10, Lemma 3.9].
Theorem 2.12
Let $E$ be a Fréchet space and $p\in[1,\infty]$. Then the map
$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$
@@
Let $p\in[1,\infty]$. Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [24, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology. Hence the sets
@@
Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map. Then $f\circ\eta\in AC_{L^{p}}([a,b],F)$ for each $\eta\in AC_{L^{p}}([a,b],V)$.
@@
Let $E$ and $F$ be sequentially complete locally convex space, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. Let $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{k+2}$-map, then the map
$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
@@
For sequentially complete locally convex space $E$ and $F$ we have
@@
For the complex case of Lemma 2.16, we will use the following result [9, Lemma 2.5].
@@
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be an open subset and $f:U\to F$ be a map. We say that $f$ is real analytic if it extends to a complex analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$.
@@
$f^{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
@@
First we consider the case $\mathbb{K}=\mathbb{C}$. By Lemma 2.16 the map $f^{*}$ is smooth and $d(f^{*})$ is complex linear in the second variable, hence by Lemma 2.18 the map $f$ is complex analytic.
If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})^{*}$ is the complex analytic extension of $f^{*}$, whence $f^{*}$ is real analytic. ∎
@@
Let $N$ be a smooth manifold modeled with on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verify
@@
For the case of absolute continuous functions with valued in a manifold, we have the following results (see [24]).
@@ 3 The space of absolutely continuous sections
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its vector bundle and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$
@@
For a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$. Since $\eta([t_{i-1},t_{i}])\subseteq U_{i}$, we have

### v4

@@ Abstract
For $p\in[1,\infty]$, we define a smooth manifold structure on the set $AC_{L^{p}}([a,b],N)$ of absolutely continuous functions $\gamma\colon[a,b]\to N$ with $L^{p}$-derivatives for all real numbers $a<b$ and each smooth manifold $N$ modeled on a sequentially complete locally convex topological vector space, such that $N$ admits a local addition. Smoothness of natural mappings between spaces of absolutely continuous functions is discussed, like superposition operators $AC_{L^{p}}([a,b],N_{1})\to AC_{L^{p}}([a,b],N_{2})$, $\eta\mapsto f\circ\eta$, for a smooth map $f\colon N_{1}\to N_{2}$. Let $G$ be a right half-Lie group. For $1\leq p<\infty$, we show certains properties of $L^{p}$-semiregularity of $G$ under certains condicions on $G$. For $r\in\mathbb{N}$, we prove that the right half-Lie groups $\text{Diff}_{K}^{r}(\mathbb{R})$ and $\text{Diff}^{r}(M)$ are $L^{p}$-semiregular. Here $K$ is a compact subset of $\mathbb{R}$ and $M$ is a compact smooth manifold (possible with boundary). An $L^{p}$-semiregular half-Lie group $G$ admits an evolution map $\text{Evol}:L^{p}([0,1],T_{e}G)\to AC_{L^{p}}([0,1],G)$, where $e$ is the neutral element of $G$. For the preceding examples, the evolution map Evol is continuous.
@@ 1 Introduction
The study of mapping manifolds is an important branch of geometry and global analysis. For example, for $k\in\mathbb{N}\cup\{0\}$, a well-known case is the manifold of the set of $C^{k}$-maps $f:M\to N$, denoted by $C^{k}(M,N)$, for a compact manifold $M$ and a smooth manifold $N$ admitting local addition (see e.g. [1, 9, 28]). In [24], Krikorian shows how to contruct smooth manifold structures for set of functions with Hölder conditions and Sobolev functions.
Our main focus will be the study of absolutely continuous functions with values in a manifold $N$ modeled on a sequentially complete locally convex space which admits a local addition. For a Hilbert manifold $N$, a smooth manifold structure on the set $AC_{L^{2}}([0,1],N)$ of absolutely continuous paths with $L^{2}$-derivatives in local charts was used by Flaschel and Klingenberg to study closed geodesics in Riemannian manifolds (see [10] and [22], cf. also [39]). For Banach manifolds $N$ admitting a smooth local addition and $p\in[1,\infty]$, Pierron and Trouvé show that a smooth manifold structure on $AC_{L^{p}}([0,1],N)$ may also be obtained using a method of Krikoian [24], which is similar to Palais’ use of Banach section functors [35]. A Lie group structure (and hence a smooth manifold structure) on $AC_{L^{p}}([0,1],G)$ was obtained in [33] for each Lie group $G$ modeled on a sequentially complete locally convex space (as in [30]), generalizing the case of Fréchet–Lie groups treated in [14].
@@
defined on an open set $\Omega\subseteq TN$ which contains the zero-vector $0_{p}\in T_{p}N$ for each $p\in N$, such that $\Sigma(0_{p})=p$ for all $p\in N$ and
@@
has open image and is a $C^{\infty}$-diffeomorphism onto its image.
Throughout the following, $a<b$ are real numbers and $p\in[1,\infty]$. For a definition of absolutely continuous functions with values in a sequentially complete locally convex space $E$ or a smooth manifold $N$ defined thereon, the reader is referred to Definitions 2.12 and 2.23, respectively (see also [33]). For $\eta\in AC_{L^{p}}([a,b],N)$, the pointwise operations make
@@
${\mathcal{V}}_{\eta}:=\{\tau\in\Gamma_{AC}(\eta)\colon\tau([a,b])\subseteq\Omega\}$
@@
${\mathcal{U}}_{\eta}:=\{\gamma\in AC_{L^{p}}([a,b],N)\colon(\eta(t),\gamma(t))\in\Omega^{\prime}\;\mbox{for all $t\in[a,b]$}\},$
@@
If, in the situation of Definition 1.1, $N$ is a $\mathbb{K}$-analytic manifold modeled on a locally convex topological $\mathbb{K}$-vector space over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\theta\colon\Omega\to\Omega^{\prime}$ is a diffeomorphism of $\mathbb{K}$-analytic manifolds, then $\Sigma$ is called a $\mathbb{K}$-analytic local addition. In this case, $AC_{L^{p}}([a,b],N)$ can be given a $\mathbb{K}$-analytic manifold structure modeled on the locally convex topological $\mathbb{K}$-vector spaces $\Gamma_{AC}(\eta)$ with properties as in Theorem 1.2, replacing $C^{\infty}$-diffeomorphisms with diffeomorphisms of $\mathbb{K}$-analytic manifolds there (see Corollary 3.6). If $N_{j}$ is a $\mathbb{K}$-analytic manifold modeled on a sequentially complete locally convex topological $\mathbb{K}$-vector space such that $N_{j}$ admits a $\mathbb{K}$-analytic local addition for $j\in\{1,2\}$, then the map $AC_{L^{p}}([a,b],f)$ described in Theorem 1.3 is $\mathbb{K}$-analytic for all $p\in[1,\infty]$ (see Corollary 3.9).
If $\pi_{TN}:TN\to N$ denotes the tangent bundle of $N$. Followings results on more general manifolds of mappings (see e.g. [3]), we can conclude that the map
$AC_{L^{p}}([a,b],\pi_{TN}):AC_{L^{p}}([a,b],TN)\to AC_{L^{p}}([a,b],N),\quad\tau\mapsto\pi_{TN}\circ\tau$
defines a vector bundle on $AC_{L^{p}}([a,b],N)$. The proofs are analogous to the reference given, but we will show them for the sake of completeness.
@@
is a smooth map and a right action of $G$ on $TG$. The following concept strengthens “regularity” in the sense of Milnor [30].
@@
If $G$ is modeled on a Fréchet space, (1.1) is required to hold for almost all $t\in[0,1]$ with respect to Lebesgue measure. In the general case, $\eta$ is required to be a Carathéodory solution to (1.1), i.e., it solves the corresponding integral equation piecewise in local charts. We mention that a Lie group $G$ is $L^{p}$-regular if and only if $G$ is $L^{p}$-semiregular and $\operatorname{Evol}$ is smooth as map
@@
(see [33]). The latter holds whenever $\operatorname{Evol}\colon L^{p}([0,1],\mathfrak{g})\to C([0,1],G)$ is continuous at 00 (see [15, Theorem E]).
Now consider a half-Lie group $G$ modeled on a sequentially complete locally convex space $E$. Thus $G$ is a group, endowed with a smooth manifold structure modeled on $E$ which makes $G$ a topological group and turns the right translation $\rho_{g}\colon G\to G$ into a smooth mapping for each $g\in G$ (cf. [6, 27]). Let us use notation as in the case of Lie groups.
@@
$\dot{y}(t)=\gamma(t).y(t),\quad t\in[0,1]$ (1.3)
satisfies local uniqueness of Carathéodory solutions for each $\gamma\in{\mathcal{L}}^{p}([0,1],\mathfrak{g})$ (in the sense of [18]) and the initial value problem (1.1) has a Carathéodory solution $\operatorname{Evol}(\gamma):=\eta\colon[0,1]\to G$.
The Lie group $\text{Diff}(M)$ of $C^{\infty}$-diffeomorphisms of a compact smooth manifold $M$ without boundary is known to be $L^{1}$-regular, and also the Lie group $\text{Diff}_{K}(\mathbb{R}^{n})$ of all $C^{\infty}$-diffeomorphisms $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that $\phi(x)=x$ for all $x\in\mathbb{R}^{n}\setminus K$, for each compact subset $K\subseteq\mathbb{R}^{n}$ (see [14]). For each positive integer $r$, the following analogues are obtained (see Theorems  6.5 and 5.3):
@@
If we replace $L^{p}$ with $L_{rc}^{\infty}$ (see Definition 2.5) the preceding theorem remains valid.
@@
Let $G$ be a right half-Lie group modeled in a sequentially complete locally convex space space $E$ which admits a local addition and $1\leq p<\infty$. Let $G$ be $L^{p}$-semiregular with continuous evolution map
@@
is continuous. If $G$ is a right half-Lie group modeled in an integral complete locally convex space $E$, then if we replace $L^{p}$ with $L_{rc}^{\infty}$ the result remains valid.
@@ 2 Preliminaries
exist for all $i,j\in\mathbb{N}\cup\{0\}$ such that $i\leq r$ and $j\leq s$, and all $y_{1},...,y_{i}\in E_{1}$ and $b_{1},...,b_{j}\in E_{2}$, and, we assume that the mappings
@@
are continuous, then $f$ is called a $C^{r,s}$-map.
@@
$\mu:\mathcal{B}(X)\to[0,\infty]$ on its $\sigma$-algebra of Borel sets and let $Y$ be a topological space. A function $\gamma:X\to Y$ is called Lusin $\mu$-measurable (or $\mu$-measurable) if for each compact subset $K\subseteq X$, there exists a sequence $(K_{n})_{n\in\mathbb{N}}$ of compact subsets $K_{n}\subseteq K$ such that each restriction $\gamma|_{K_{n}}$ is continuous and $\mu\left(K\backslash\cup_{n\in\mathbb{N}}K_{n}\right)=0$.
@@
Let $E$ be a locally convex space, $a<b$ be real numbers, $1\leq p<\infty$ and $\lambda:\mathcal{B}([a,b])\to[0,\infty)$ be the restriction of the Lebesgue-Borel measure on $\mathbb{R}$. We define the set $\mathcal{L}^{p}([a,b],E)$ as the set of all Lusin $\lambda$-measurable functions $\gamma:[a,b]\to E$ such that for each continuous seminorm $q$ on $E$ we have
@@
Let $\gamma\sim\eta$ if and only if $\gamma(t)=\eta(t)$ for almost all $t\in[a,b]$ and write $[\gamma]$ for the equivalence class of $\gamma$. We define the Hausdorff locally convex space
@@
$\lVert\gamma\lVert_{\mathcal{L}^{\infty},q}:=\operatorname*{ess\,sup}_{t\in[a,b]}q\circ\gamma(t).$
@@
We define the vector space $\mathcal{L}_{rc}^{\infty}([a,b],E)$ of all Borel measurable functions $\gamma:[a,b]\to E$ such that the image $\text{Im}(\gamma)$ has compact and metrizable closure, endowed with the topology induced by $\mathcal{L}^{\infty}([a,b],E)$. Thus
@@
Let $E$ be a Frechet space and $p\in[1,\infty]$. If $L_{\mathcal{B}}^{p}([0,1],E)$ denotes the Lebesgue space constructed with the set of Borel measurable functions (see [14]), then $L_{\mathcal{B}}^{p}([0,1],E)$ coincides with $L^{p}([0,1],E)$ ([33, Proposition 2.10]).
@@
Let $E$ be a locally convex space and ${\gamma:[a,b]\to E}$ be such that $\alpha\circ\gamma\in\mathcal{L}^{1}([a,b],\mathbb{R})$ for each $\alpha\in E^{\prime}$. An element $z\in E$ is called the weak integral of $\gamma$ if
@@
for each $\forall\alpha\in E^{\prime}$. Then $z$ is called the weak integral of $\gamma$ from $a$ to $b$, and we write $z=:\int_{a}^{b}\gamma(s)ds$.
Definition 2.9
Let $E$ be a locally convex space. We say that a sequence $(x_{n})_{n}\subset E$ is a Cauchy sequence if for each $\varepsilon>0$ and each continuous seminorm $q$ of $E$, there exists an integer $N_{\varepsilon,q}\in\mathbb{N}$, such that for all $m,n\geq N_{\varepsilon,q}$ we have
$q(x_{m}-x_{n})<\varepsilon.$
We say that $E$ is sequentially complete if every Cauchy sequence converge in $E$.
The following lemma [33, Proposition 2.26] allows us to define absolute continuous functions with vector values.
@@
We denote the space of all $L^{p}$-absolutely continuous functions by $AC_{L^{p}}([a,b],E)$. Let $t_{0}\in[a,b]$ be fixed, since $\eta^{\prime}:=[\gamma]$ is necessarily unique (see [32, Lemma 2.28]), the map
$\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E),\quad\eta\mapsto(\eta(t_{0}),\eta^{\prime})$ (2.2)
is an isomorphism of vector spaces. We endow $AC_{L^{p}}([a,b],E)$ with the Hausdorff locally convex vector topology which makes $\Phi$ an isomorphism of topological vector space (see [32, Definition 3.1]).
The following result will allow us to study differentiability of functions with values in $AC_{L^{p}}([a,b],E)$.
Lemma 2.13
Let $E$ be a sequentially complete locally convex space and $p\in[1,\infty]$. Then the map
$\Psi:AC_{L^{p}}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad\eta\mapsto(\eta,\eta^{\prime})$ (2.3)
@@
Proof.
We let $I:E\times L^{p}([a,b],E)\to C([a,b],E)$ be the continuous map given by
$I((x,[\gamma])(t):=x+\int_{a}^{t}\gamma(s)ds,\quad t\in[a,b]$
for each $x\in E$ and $[\gamma]\in L^{p}([a,b],E)$. Let $\Phi:AC_{L^{p}}([a,b],E)\to E\times L^{p}([a,b],E)$, with $t_{0}:=a$, be the isomorphism of topological vector spaces as above. We consider the map
$\Theta:E\times L^{p}([a,b],E)\to C([a,b],E)\times{L^{p}}([a,b],E),\quad(x,[\gamma])\mapsto\big{(}I(x,[\gamma]),[\gamma]\big{)}$
which is continuous.
Moreover, since the evaluation map $\varepsilon_{a}:C([a,b],E)\to E$, $\eta\mapsto\eta(a)$ is continuous, the map $\left(\Theta|^{\text{Im}(\Theta)}\right)^{-1}=(\varepsilon_{a},\text{id}_{L^{p}})$ is also continuous. Hence $\Psi$ is a topological embedding.
Let $(\eta_{\alpha},\eta_{\alpha}^{\prime})_{\alpha}$ be a net in $\text{Im}(\Theta)$ that converges to $(\eta,[\gamma])\in C([a,b],E)\times{L^{p}}([a,b],E)$. By continuity of $\varepsilon_{a}$, the net $(\eta_{\alpha}(a))_{\alpha}$ converges to $\eta(a)\in E$, and by continuity of $\Theta$, the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ converges to $\left(I(\eta(a),[\gamma]),[\gamma]\right)\in\text{Im}(\Theta)$. Since the net $\left(\Theta(\eta_{\alpha}(a),\eta_{\alpha}^{\prime})\right)_{\alpha}$ also converges to $(\eta,[\gamma])$, we have
$I(\eta(a),[\gamma])=\eta.$
Therefore $\eta^{\prime}=[\gamma]$ and $(\eta,[\gamma])\in\text{Im}(\Theta)$. ∎
@@
Let $p\in[1,\infty]$. Since the inclusion map ${AC_{L^{p}}([a,b],E)\to C([a,b],E)}$ is continuous [33, Lemma 3.2], the topology on $AC_{L^{p}}([a,b],E)$ is independent of the choice of $t_{0}$ and finer than the compact-open topology. Hence the sets
@@
Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$, $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{1}$-map. Then $f\circ\eta\in AC_{L^{p}}([a,b],F)$ for each $\eta\in AC_{L^{p}}([a,b],V)$.
@@
Let $E$ and $F$ be sequentially complete locally convex spaces, $p\in[1,\infty]$ and $k\in\mathbb{N}\cup\{0,\infty\}$. Let $V\subseteq E$ be open subset and ${f:V\to F}$ be a $C^{k+2}$-map, then the map
$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
@@
For sequentially complete locally convex spaces $E$ and $F$ we have
@@
Definition 2.19
Let $E$ and $F$ be complex topological vector spaces, where is $F$ is locally convex, $U\subseteq E$ be an open subset and $f:U\to F$ be a mapping. We say that $f$ is complex analytic if it is continuous and, for each $x\in U$, there exists a 00-neighbouhood $V\subseteq E$ such that $x+V\subseteq U$ and certain continuous homogeneous polynomials $\beta_{n}:E\to F$ of degree $n$, such that $f$ admits the expansion: $f(x+y)=\sum_{n=0}^{\infty}\beta_{n}(y)$, for all $y\in V$.
For our context, we present an application of [4, Proposition 7.7] to our particular case.
@@
Let $E$ and $F$ be real locally convex spaces, $U\subseteq E$ be an open subset and $f:U\to F$ be a map. We say that $f$ is real analytic if it extends to an analytic map $f:V\to F_{\mathbb{C}}$ on some open neighborhood $V$ of $U$ in $E_{\mathbb{C}}$, where $E_{\mathbb{C}}$ and $F_{\mathbb{C}}$ denotes the complexification of $E$ and $F$, respectively.
@@
$f_{*}:=AC_{L^{p}}([a,b],f):AC_{L^{p}}([a,b],V)\to AC_{L^{p}}([a,b],F),\quad\eta\mapsto f\circ\eta$
@@
First we consider the case $\mathbb{K}=\mathbb{C}$. By Lemma 2.17 the map $f_{*}$ is smooth and $d(f_{*})$ is complex linear in the second variable, hence by Lemma 2.20 the map $f_{*}$ is complex analytic.
If $\mathbb{K}=\mathbb{R}$, then by definition the map $f$ has a complex analytic extension $\tilde{f}$, hence $(\tilde{f})_{*}$ is the complex analytic extension of $f_{*}$, whence $f_{*}$ is real analytic. ∎
@@
Let $N$ be a smooth manifold modeled with on a sequentially complete locally convex space $E$ and $p\in[1,\infty]$. We say that a function $\eta:[a,b]\to N$ is $L^{p}$-absolutely continuous if it is continuous and there exists a partition $\{t_{0},...,t_{n}\}$ of $[a,b]$ such that for each $j\in\{1,...,n\}$, there exists a chart $\varphi_{j}:U_{j}\to V_{j}$ that verifies
@@
For the case of absolutely continuous functions with values in a manifold, the following facts are available (see [33]).
@@
Let $N$ be a smooth manifold modeled on a sequentially complete locally convex space $E$, $\pi_{TN}:TN\to N$ its tangent bundle and $p\in[1,\infty]$. For $\eta\in AC_{L^{p}}([a,b],N)$ we define the set
${\Gamma_{AC}(\eta)}:=\{\sigma\in AC_{L^{p}}([a,b],TN):\pi_{TN}\circ\sigma=\eta\}$ (2.4)
@@
Consider a partition $P_{n}=\{t_{0},...,t_{n}\}$ of $[a,b]$ and charts $\{(\varphi_{i},U_{i}):i\in\{1,..,n\}\}$ of $N$ that verify the definition of absolute continuity for $\eta$. Since $\eta([t_{i-1},t_{i}])\subseteq U_{i}$, we have